((-10)/(-10)-3)/(-2)
Answer by unlockmath(1688)

Hello,
With this ((-10)/(-10)-3)/(-2) we first handle:
((-10)/(-10)-3)=-1/3
Now the next is:
-1/3/(-2)= 2/3
Make sense:
RJ
www.math-unlock.com